A GO bus is travelling along the QEW when it enters a construction zone. This causes the GO bus to decelerate at 1.80 m/s . If it takes the GO bus 35 s to stop, calculate the displacement mtravelled by the GO bus to stop? 


1
Expert's answer
2021-01-15T09:33:10-0500

The displacement is


"d=\frac{at^2}{2}=\frac{1.8\cdot35^2}{2}=1102.5\text{ m}."